The perimeter of a rectangle is 42 ft. The length of the rectangle 3 ft. less than twice the width.Find the length and width of the rectangle.
Answer by richard1234(7193)

If x and 2x - 3 are the width and length of the rectangle, then
2(x + 2x-3) = 42
x + (2x - 3) = 21
Solving, we obtain x = 8, 2x - 3 = 13.